# On Hardness and Approximation of Broadcasting in Sparse Graphs

### v1

@@ Abstract
We study the Telephone Broadcasting problem in sparse graphs. Given a designated source in an undirected graph, the task is to disseminate a message to all vertices in the minimum number of rounds, where in each round every informed vertex may inform at most one uninformed neighbor. For general graphs with $n$ vertices, the problem is NP-hard. Recent work shows that the problem remains NP-hard even on restricted graph classes such as cactus graphs of pathwidth $2$ [Aminian et al., ICALP 2025] and graphs at distance-1 to a path forest [Egami et al., MFCS 2025].
In this work, we investigate the problem in several sparse graph families. We first prove NP-hardness for $k$-cycle graphs, namely graphs formed by $k$ cycles sharing a single vertex, as well as $k$-path graphs, namely graphs formed by $k$ paths with shared endpoints. Despite multiple efforts to understand the problem in these simple graph families, the computational complexity of the problem had remained unsettled, and our hardness results answer open questions by Bhabak and Harutyunyan [CALDAM 2015] and Harutyunyan and Hovhannisyan [COCAO 2023] concerning the problem’s complexity in $k$-cycle and $k$-path graphs, respectively.
On the positive side, we present Polynomial-Time Approximation Schemes (PTASs) for $k$-cycle and $k$-path graphs, improving over the best existing approximation factors of $2$ for $k$-cycle graphs and an approximation factor of $4$ for $k$-path graphs. Moreover, we identify a structural frontier for tractability by showing that the problem is solvable in polynomial time on graphs of bounded bandwidth. This result generalizes existing tractability results for special sparse families such as necklace graphs.
@@ 1 Introduction
The problem is known to be computationally difficult. Slater et al.  [19] proved that Telephone Broadcasting is NP-complete. Elkin and Kortsarz [7] established that approximating Telephone Broadcasting within a factor of $3-\epsilon$ for any $\epsilon>0$ is NP-hard. On the other hand, Elkin and Kortsarz [8] designed an $\mathcal{O}(\log n/\log\log n)$-approximation algorithm for general graphs, which is the best existing approximation algorithm. Whether a constant-factor approximation exists for general graphs remains open, though such results are known for specific classes, including unit disk graphs [18] and cactus graphs [1].
A substantial body of work has investigated the complexity of Telephone Broadcasting on restricted graph classes. On the positive side, polynomial-time algorithms are known for several families, including trees [10], grids, stars of cliques, fully connected trees, chains of rings, necklace graphs, and certain subclasses of cactus graphs; see, e.g., [5, 12, 17, 20, 9]. On the negative side, hardness results persist even for structurally simple graphs: the problem is NP-hard in planar cubic graphs and in graphs of bounded treewidth. More recently, Tale showed NP-hardness for graphs of pathwidth $3$, while Aminian et al. [1] established hardness for so-called snowflake graphs (cactus graphs of pathwidth $2$), resolving a long-standing open question about the complexity of broadcasting in cacti. Even stronger results were obtained by Egami et al. [6], who proved NP-hardness for graphs at distance 1 from path forests, i.e., graphs in which deleting a single vertex yields a path forest. See Figure 1 for illustrations of snowflake graphs and graphs at distance 1 from path forests.
These results highlight intriguing frontier cases where the complexity of the problem remains unresolved. One such case is that of $k$-cycle graphs (also called flower graphs), obtained by joining $k$ cycles at a single vertex. Despite their simple definition, these graphs exhibit surprisingly rich behavior. Broadcasting on $k$-cycle graphs has been studied from an approximation perspective: the best known algorithm achieves a factor of 2 [2], yet the exact complexity remains open. A closely related family is that of $k$-path graphs, formed by joining $k$ paths at their endpoints. Several approximation algorithms are known for $k$-path graphs [3, 14], with the current best result being a 2-approximation [14]. Nevertheless, the precise computational complexity of the problem for this family also remains open.
@@ 1.1 Contribution
  • We prove that Telephone Broadcasting is NP-Complete in two of the simplest sparse graph families: $k$-cycle graphs (Theorem 3.4) and $k$-path graphs (Theorem 3.5). Both families have been studied in the broadcasting literature, yet their complexity status remained open. Our hardness results also imply hardness for broader classes, including series–parallel graphs (SP-graphs), for which no hardness was previously known.
  • We design polynomial-time approximation schemes (PTAS) for both $k$-cycle graphs (Corollary 4.7) and $k$-path graphs (Corollary 4.14). To the best of our knowledge, these are the first graph families for which Telephone Broadcasting admits both a hardness result and a matching PTAS, highlighting their role as boundary cases between tractability and intractability.
  • We present a polynomial-time algorithm for optimal broadcasting in graphs of bounded bandwidth (Theorem 5.2). This result generalizes several known tractable cases, such as chains of rings and related structures. Notably, Telephone Broadcasting is among the few problems exhibiting a complexity gap between graphs of bounded bandwidth (where the problem is tractable) and graphs of bounded pathwidth (where the problem is NP-hard). In particular, $k$-cycle and $k$-path graphs both have pathwidth 2 but unbounded bandwidth, placing them precisely on the intractable side of this gap.
(a) A snowflake graph (cactus of pathwdith 2) [1]
(b) Graph of distance 1 to path forest [6]
(c) A $k$-cycle graph (Theorem 3.4)
(d) A $k$-path graph (Theorem 3.5)
@@ 1.2 Overview of Techniques
Our results combine hardness reductions, approximation schemes, and dynamic programming, each tailored to the structural properties of the graph families under consideration. We give a brief outline here.
@@
  • For the negative results, we design reductions that encode scheduling constraints directly into the graph structure. In particular, the NP-hardness proofs for $k$-cycle and $k$-path graphs exploit the interaction between cycle (or path) lengths and the timing of broadcast calls. The shared vertex in $k$-cycles (or the pair of endpoints in $k$-paths) acts as the originator, and we show that deciding feasibility becomes equivalent to a combinatorial partitioning problem whose complexity propagates to the broadcasting instance. Our reductions build a more restricted variant of Numerical 3-Dimensional Matching (RN3DM) (Definition 3.1), previously employed in hardness proofs for telephone broadcasting. This restricted variant, which is known to be still NP-hard [22], makes it possible to encode number-theoretic constraints using only these simple graph families.
@@
  • We design polynomial-time approximation schemes for $k$-cycle and $k$-path graphs by reducing the broadcasting problem to covering problems over integers. For $k$-cycle graphs, the task reduces to covering a multiset of cycle lengths $\{\ell_{1},\ldots,\ell_{k}\}$ with the smallest possible broadcast round $\beta$ using pairs of numbers from $\beta\,]$, so that for each $\ell_{i}$ there exist $a,b\in[\,\beta\,]$ associated with $\ell_{i}$ such that $a+b\geq\ell_{i}$ (See Definition 4.1). For $k$-path graphs, the covering problem is slightly different: we seek the smallest $\beta$ such that each $\ell_{i}$ can be covered by one number from $\beta\,]$ and one from $\beta-d\,]$, where $d$ is the length of the shortest path (see Definition 4.9). In both cases, approximate solutions to the number-theoretic covering problems translate into near-optimal broadcasting schedules, yielding PTASs for $k$-cycle and $k$-path graphs.
@@
  • For finding optimal broadcast schemes in graphs of bounded bandwidth, we exploit the fact that a bandwidth-$k$ ordering restricts each vertex to neighbors within a sliding window of $\mathcal{O}(k)$ vertices. This allows us to describe the status of any partial broadcast tree using a compact boundary state and use a dynamic programming approach which has a table with $n$ rows (one for each prefix of the ordering) and at most $n^{\mathcal{O}(k^{2})}$ states per row, giving a total size bounded by $f(k)\cdot n^{\mathcal{O}(k^{2})}$.
@@ 2 Preliminaries
We use notation $[n]$ and $[n,m]$ for positive integers $n$ and $m$ for the set of all natural numbers less than or equal to $n$ and all natural numbers between $n$ and $m$ (inclusive), respectively (i.e. $n]=\{1,2,\ldots,n\}$ and $n,m]=\{n,n+1,\ldots,m\}$). Also, we refer to the number of distinct elements in a multiset as its “support”.
@@
Definition 2.2 ($k$-cycle and $k$-path graphs).
A $k$-cycle graph $G$ contains $k$ cycles with exactly one common vertex $r$. A $k$-path graph $G$ contains $k$ paths with shared endpoints $s$ and $t$.
See Figure 1 (parts (c) and (d)) for illustrations of $k$-cycle and $k$-path graphs. Observe that any $k$-cycle (respectively, $k$-path) graph can be fully described by a set of integers specifying the number of vertices on each cycle (respectively, path). This representation allows us to relate such graphs directly to integer sets, a connection that underlies both our hardness proofs and our algorithmic techniques. These graph families are also simple representatives of well-studied classes: both have bandwidth 2, $k$-cycle graphs form a subclass of cactus graphs, and $k$-path graphs form a subclass of series–parallel graphs. Consequently, our hardness results extend beyond $k$-cycles and $k$-paths to broader sparse families, including series–parallel graphs, for which no hardness results were previously known.
Definition 2.3 (Graph Bandwidth).
Given a graph $G$, an ordering with bandwidth $k$ of $G$ is a permutation $v_{1},\ldots,v_{n}$ of vertices of $G$ such that for each edge $v_{i}\sim v_{j}$ we have $|i-j|\leq k$. The bandwidth of $G$ is the minimum $k$ such that such an ordering exists.
Certain graphs of bounded bandwidth have also been studied in the context of telephone broadcasting. For instance, necklace graphs—collections of cycles where each consecutive pair shares a vertex—have bandwidth $k=2$ and the problem can be optimally solved for them [13, 15].
@@ 3 Hardness Results
In this section, we present our hardness results for Telephone Broadcasting in $k$-cycle graphs and $k$-path graphs. These results extend the existing hardness landscape by showing that even some of the simplest graph families exhibit computational intractability.
Recent hardness results for broadcasting typically rely on reductions from variants of the Numerical Matching with Target Sums (NMTS) problem, which is known to be strongly NP-hard [11] (Problem SP17). An instance of NMTS consists of three multisets $U,V,$ and $W$, each containing $m$ positive integers. The decision problem asks whether there exists a partition of $U\cup V\cup W$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, such that each $A_{i}$ contains exactly one element $u_{i}\in U$, one element $v_{i}\in V$, and one element $w_{i}\in W$, with the property that $u_{i}+v_{i}=w_{i}$ for all $i\in[m]$.
Tale [21] introduced a variant of NMTS, called Numerical 3-Dimensional (Almost) Matching, in which no repeated entries appear in $U$ and $V$, and all elements of $U\cup V\cup W$ are multiples of $m$ and two integers $T$ and $V$. As in NMTS, the objective is to partition $U\cup V$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, each containing exactly one element from $U$ and one element from $V$. However, instead of requiring $\sum_{a\in A_{i}}a=w_{i}$, the condition is relaxed to $\sum_{a\in A_{i}}a\geq T-\lambda.$ Tale showed that this problem is strongly NP-hard via a reduction from NMTS, and subsequently reduced it to Telephone Broadcasting in graphs of pathwidth at most 3. Egami et al. [6] studied another restricted variant of NMTS, in which all members of $U$ are even while all members of $V\cup W$ are odd. They proved that this variant remains strongly NP-hard and provided a reduction to broadcasting in graphs that are distance-1 from a path forest.
@@
Given a multiset $W=\{w_{1},w_{2},\ldots,w_{m}\}$ of positive integers such that $\sum_{i=1}^{m}w_{i}+m(m+1)=me$ for some integer $e$, do there exist two permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)+w_{i}=e\quad\text{for all }1\leq i\leq m\,$?
@@
Consider the instance of RN3DM with $W=\{1,3,4,4\}$, so $m=4$. Note that we have $\sum_{i=1}^{m}w_{i}+m(m+1)=32=me$ for $e=8$. Let $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$. One can check that for every $i\in[4]$, $\lambda(i)+\mu(i)+w_{i}=e.$ For instance, $\lambda(1)+\mu(1)+w_{1}=4+3+1=8$, and $\lambda(2)+\mu(2)+w_{2}=1+4+3=8.$ The same holds for $i=3,4$. Hence $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
Given that RN3DM is strongly NP-hard [22], we assume that the numbers in $W$ are encoded in unary. We present reductions from RN3DM to broadcasting in $k$-cycle and $k$-path graphs. For $k$-cycle graphs, our approach introduces an intermediate problem as part of the reduction series, whereas for $k$-path graphs, we directly reduce RN3DM to telephone broadcasting.
3.1 Hardness of Telephone Broadcasting in $k$-cycle graphs
We prove that the $k$-cycle broadcasting decision problem is NP-hard through a series of reductions starting from RN3DM. As a first step, we define an intermediate problem, called Even-Odd RN3DM, which is a variant of RN3DM. In this variant, instead of using two permutations $\lambda$ and $\mu$ of $[m]$ (matched with the input set $W$), we use $\alpha$ as a permutation of the even integers in $[2m]$ and $\beta$ as a permutation of the odd integers in $[2m]$ to be matched with an input set that we call $C$. Formally, Even-Odd RN3DM is defined as follows.

### v2

# On Hardness and Approximation of Broadcasting in Structured Graphs

@@ Abstract
We study the Telephone Broadcasting problem in graphs with restricted structure. Given a designated source in an undirected graph, the goal is to disseminate a message to all vertices in the minimum number of rounds, where in each round every informed vertex may inform at most one uninformed neighbor. For general graphs with $n$ vertices, the problem is NP-hard. Recent work shows that the problem remains NP-hard even on restricted graph classes such as graphs of treewidth 2 [Tale 2025], cactus graphs of pathwidth $2$ [Aminian et al., ICALP 2025] and graphs at distance-1 to a path forest [Egami et al., MFCS 2025].
In this work, we investigate the problem in several sparse graph families. We first prove NP-hardness for cycle-star graphs, namely graphs formed by $k$ cycles sharing a single vertex, as well as melon graphs, namely graphs formed by $k$ paths with shared endpoints. Despite multiple efforts to understand the problem in these simple graph families, the computational complexity of the problem remained unsettled. Our hardness results answer open questions by Bhabak and Harutyunyan [CALDAM 2015] and Harutyunyan and Hovhannisyan [COCOA 2023] concerning the problem’s complexity in cycle-star and melon graphs, respectively.
On the positive side, we present Efficient Polynomial-Time Approximation Schemes (EPTASs) for cycle-star and melon graphs, improving over the best existing approximation factors of $2$ for both graph families. Moreover, we identify a structural frontier for tractability by showing that the problem is solvable in polynomial time on graphs of bounded cutwidth, a class that generalizes other families such as graphs of bounded bandwidth. This result subsumes existing tractability results for special sparse graph families such as necklace graphs.
Finally, for split graphs, a fundamental class of highly structured graphs, we obtain a polynomial-time algorithm with approximation factor 1.76. This improves on the previously known factor $2$ bound; the same approach also applies to the multi-source setting.
@@ 1 Introduction
The problem is known to be computationally difficult. Slater et al.  [25] proved that Telephone Broadcasting is NP-Complete. Elkin and Kortsarz [9] established that approximating Telephone Broadcasting within a factor of $3-\epsilon$ for any $\epsilon>0$ is NP-hard. On the other hand, Elkin and Kortsarz [10] designed an $\mathcal{O}(\log n/\log\log n)$-approximation algorithm for general graphs, which is the best existing approximation algorithm. Whether a constant-factor approximation exists for general graphs remains open, though such results are known for specific classes, including unit disk graphs [24], cactus graphs [2], and graphs of bounded pathwidth [2, 8].
A substantial body of work has investigated the complexity of Telephone Broadcasting on restricted graph classes. On the positive side, polynomial-time algorithms are known for several families, including trees [12], grids, stars of cliques, fully connected trees, necklace graphs (chains of rings), and certain subclasses of cactus graphs; see, e.g., [7, 14, 23, 26, 11]. On the negative side, hardness results persist even for structurally simple graphs such as planar cubic graphs [19]. More recently, Tale [27] showed that the problem remains NP-hard on graphs at distance 2 from path forests, as well as graphs of pathwidth 3 and treewidth 2. Aminian et al. [2] established hardness for so-called snowflake graphs (cactus graphs of pathwidth $2$). Even stronger results were obtained by Egami et al. [8], who proved NP-hardness for graphs at distance 1 from path forests, i.e., graphs in which deleting a single vertex yields a path forest. Figure 1 illustrates several hard sparse graph instances.
These results highlight intriguing frontier cases where the problem’s complexity remains unresolved. One such case is that of cycle-star graphs (also called flower graphs), obtained by joining $k$ cycles at a single vertex. Despite their simple definition, these graphs exhibit surprisingly rich behavior. Telephone Broadcasting on cycle-star graphs has been studied from an approximation perspective: the best known algorithm achieves a factor of 2 [3], yet the exact complexity remains open. A closely related family is that of melon graphs, formed by joining $k$ paths at their endpoints. Several approximation algorithms are known for melon graphs [4, 15], with the current best result being a 2-approximation [15]. Nevertheless, the precise computational complexity of the problem for this family also remains open. Another structurally interesting graph family is that of split graphs, which consist of a clique and an independent set and therefore fall outside the sparse regime. For single-source Telephone Broadcasting, a 2-approximation algorithm is known for split graphs [16]. Moreover, the multi-source variant of Telephone Broadcasting is NP-hard even on split graphs [20], underscoring that hardness persists despite strong structural restrictions.
@@ 1.1 Contribution
  • We prove that Telephone Broadcasting is NP-Complete in two of the simplest sparse graph families: cycle-star graphs (Theorem 3.4) and melon graphs (Theorem 3.5). Both families have been studied in the broadcasting literature, yet their complexity status remained open. Our hardness results also imply hardness for broader classes, including series–parallel graphs (SP-graphs), for which no hardness was previously known.
  • We design efficient polynomial-time approximation schemes (EPTASs) for both cycle-star graphs (Corollary 4.7) and melon graphs (Corollary 4.14). To the best of our knowledge, these are the first known graph families for which Telephone Broadcasting admits both a hardness result and a matching EPTAS, highlighting their role as boundary cases between tractability and intractability. Our EPTASs extend naturally to the multi-source setting.
  • We present a polynomial-time algorithm for optimal broadcasting in graphs of bounded cutwidth (Theorem 5.2). This result generalizes several previously known tractable cases, such as necklace graphs (chains of rings) and related structures [17], and also applies to the multi-source setting. Since graphs of bounded bandwidth also have bounded cutwidth,  [6], our result further implies tractability for graphs of bounded bandwidth.
    Notably, Telephone Broadcasting is among the few problems exhibiting a complexity gap between graphs of bounded cutwidth (where the problem is tractable) and graphs of bounded pathwidth (where the problem is NP-hard [27]). In particular, cycle-star and melon graphs both have pathwidth 2 but unbounded cutwidth, placing them precisely on the intractable side of this gap.
  • We present a $1.76$-approximation algorithm for Telephone Broadcasting on split graphs (Theorem 6.2), improving over the known factor 2 bound [16]. Together with our tractability results for bounded-cutwidth graphs and hardness results for sparse families, this places split graphs as an intermediate structural class. The algorithm applies to both single-source and multi-source settings, the latter known as being APX-hard [20].
(a) A graph of distance 2 to path forest [27]
(b) A snowflake graph (cactus of pathwidth 2) [2]
(c) Graph of distance 1 to path forest [8]
(d) A cycle-star graph (Theorem 3.4)
(e) A melon graph (Theorem 3.5)
@@ 1.2 Overview of Techniques
Our results combine hardness reductions, approximation schemes, and dynamic programming, together with fixed-dimension integer programming and min-cost flow formulations, each tailored to the structural properties of the graph families under consideration. We give a brief outline here.
@@
  • For the negative results, we design reductions that encode scheduling constraints directly into the graph structure, adapting and extending hardness techniques introduced by Tale [27] and Egami et al. [8]. In particular, the NP-hardness proofs for cycle-star and melon graphs exploit the interaction between cycle (or path) lengths and the timing of broadcast calls. The shared vertex in cycle-star graphs (or the pair of endpoints in melon graphs) acts as the originator, and we show that deciding feasibility becomes equivalent to a combinatorial partitioning problem whose complexity propagates to the Telephone Broadcasting instance. Our reductions build a more restricted variant of Numerical 3-Dimensional Matching (RN3DM) (Definition 3.1), previously employed in hardness proofs for Telephone Broadcasting by Tale [27] and Egami et al. [8]. The restricted variant that we use, which is known to be still NP-hard [29], makes it possible to encode number-theoretic constraints using only these simple graph families.
@@
  • We design efficient polynomial-time approximation schemes for cycle-star and melon graphs by reducing the Telephone Broadcasting problem to covering problems over integers. For cycle-star graphs, the task reduces to covering a multiset of cycle lengths $\{\ell_{1},\ldots,\ell_{k}\}$ with the smallest possible broadcast round $\beta$ using pairs of numbers from $\beta\,]$, such that for each $\ell_{i}$ there exist $a,b\in[\,\beta\,]$ with $a+b\geq\ell_{i}$ (see Definition 4.1). For melon graphs, the covering problem is slightly different: we seek the smallest $\beta$ such that each $\ell_{i}$ can be covered by one number from $\beta\,]$ and one from $\beta-d\,]$, where $d$ is the length of the shortest path between two endpoints (see Definition 4.9). In both cases, approximate solutions to the resulting number-theoretic covering problems translate into near-optimal broadcasting schedules; moreover, the rounded covering instances have constant support and can be solved exactly using fixed-dimension integer linear programming, yielding EPTASs for cycle-star and melon graphs.
@@
  • For finding optimal broadcast schemes in graphs of bounded cutwidth, we exploit the fact that a cutwidth-$k$ ordering restricts each vertex to neighbors within a sliding window of $\mathcal{O}(k)$ vertices. This allows us to describe the status of any partial broadcast tree using a compact boundary state and use a dynamic programming approach which has a table with $n$ rows (one for each prefix of the ordering) and at most $f(k)\cdot n^{k}$ states per row, giving a total size bounded by $f(k)^{2}\cdot n^{2k+1}$, in which $f(.)$ is a function of $k$ that is independent of $n$.
  • For split graphs, we establish that informing independent-set vertices early is never beneficial (Lemma 6.1); equivalently, one can assume without loss of optimality that broadcasts to the independent set are deferred as long as possible. This structural insight allows us to decouple the broadcast process at each vertex into an initial phase within the clique, followed by a controlled dissemination to the independent set. As a result, the problem becomes related to constructing minimum-degree and minimum-overload coverings of the independent set by clique vertices. When discussing these covering problems, at a high level, we seek to minimize the number of clique vertices that are responsible for covering many independent-set vertices. Such high-load clique vertices may constitute the main bottleneck of the broadcast process and are therefore informed earlier, so that they can switch to informing independent-set vertices sooner than others. We compute a suitable covering via a reduction to a minimum-cost flow problem, yielding an approximation algorithm with approximation factor $1.76$ for both the single-source and multi-source settings.
@@ 2 Preliminaries
We use notation $[n]$ and $[n,m]$ for positive integers $n$ and $m$ for the set of all positive integers less than or equal to $n$ and all positive integers between $n$ and $m$ (inclusive), respectively (i.e. $n]=\{1,2,\ldots,n\}$ and $n,m]=\{n,n+1,\ldots,m\}$). Also, we refer to the number of distinct elements in a multiset as its “support”.
@@
We also consider a multi-source variant of Telephone Broadcasting, in which a set $S\subseteq V$ of vertices is initially informed at round 0. The broadcasting process proceeds as before, with each informed vertex informing at most one uninformed neighbor per round. The goal is again to minimize the number of rounds until all vertices are informed. Throughout the paper, unless explicitly stated otherwise, Telephone Broadcasting denotes the single-source broadcasting problem.
Definition 2.2 (Cycle-star and Melon Graphs).
A cycle-star graph $G$ contains $k$ cycles with exactly one common vertex $r$. A melon graph $G$ contains $k$ paths with shared endpoints $s$ and $t$.
See Figure 1 (parts (c) and (d)) for illustrations of cycle-star and melon graphs. Observe that any cycle-star (respectively, melon) graph can be fully described by a set of integers specifying the number of vertices on each cycle (respectively, path). This representation allows us to relate such graphs directly to integer sets, a connection that underlies both our hardness proofs and our algorithmic techniques. These graph families are also simple representatives of well-studied classes: cycle-star graphs form a subclass of cactus graphs, and melon graphs form a subclass of series–parallel graphs. Consequently, our hardness results extend beyond cycle-star and melon graphs to broader sparse families, including series–parallel graphs, for which no hardness results were previously known.
Definition 2.3 (Graph Cutwidth).
The cutwidth of a graph $G=(V,E)$, denoted $\mathrm{cw}(G)$, is the minimum integer $k$ such that there exists a linear ordering $\pi=(v_{1},\ldots,v_{n})$ of $V$ with the property that, for every $i\in[n]$, at most $k$ edges have one endpoint in $\{v_{1},\ldots,v_{i}\}$ and the other in $\{v_{i+1},\ldots,v_{n}\}$.
Computing the cutwidth is known to be fixed-parameter tractable. In particular, for every fixed constant $k$, there exists a linear-time algorithm that either constructs a cutwidth-$k$ vertex ordering or correctly reports that the cutwidth exceeds $k$ [28]. This makes cutwidth a particularly well-suited parameter for algorithmic approaches based on linear layouts and dynamic programming, such as those developed in this paper.
Definition 2.4 (Split graphs).
A graph $G=(V,E)$ is a split graph if its vertex set can be partitioned into two sets $C$ and $I$ such that the subgraph induced by $C$ is a clique and the subgraph induced by $I$ is an independent set.
@@ 3 Hardness Results
In this section, we present our hardness results for Telephone Broadcasting in cycle-star graphs and melon graphs. These results extend the existing hardness landscape by showing that even some of the simplest graph families exhibit computational intractability.
Recent hardness results for Telephone Broadcasting typically rely on reductions from variants of the Numerical Matching with Target Sums (NMTS) problem, which is known to be strongly NP-hard (Problem SP17 of  [13]). An instance of NMTS consists of three multisets $U,V,$ and $W$, each containing $m$ positive integers. The decision problem asks whether there exists a partition of $U\cup V\cup W$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, such that each $A_{i}$ contains exactly one element $u_{i}\in U$, one element $v_{i}\in V$, and one element $w_{i}\in W$, with the property that $u_{i}+v_{i}=w_{i}$ for all $i\in[m]$.
Tale [27] introduced a variant of NMTS, called Numerical 3-Dimensional (Almost) Matching, in which no repeated entries appear in $U$ and $V$, and all elements of $U\cup V\cup W$ are multiples of $m$ and two integers $T$ and $V$. As in NMTS, the objective is to partition $U\cup V$ into $m$ disjoint sets $A_{1},\ldots,A_{m}$, each containing exactly one element from $U$ and one element from $V$. However, instead of requiring $\sum_{a\in A_{i}}a=w_{i}$, the condition is relaxed to $\sum_{a\in A_{i}}a\geq T-\lambda.$ Tale showed that this problem is strongly NP-hard via a reduction from NMTS, and subsequently reduced it to Telephone Broadcasting in graphs of pathwidth at most 3. Egami et al. [8] studied another restricted variant of NMTS, in which all members of $U$ are even while all members of $V\cup W$ are odd. They proved that this variant remains strongly NP-hard and provided a reduction to Telephone Broadcasting in graphs that are distance-1 from a path forest.
@@
Given a multiset $W=\{w_{1},w_{2},\ldots,w_{m}\}$ of positive integers such that $\sum_{i=1}^{m}w_{i}+m(m+1)=m\cdot e$ for some integer $e$, do there exist two permutations $\lambda$ and $\mu$ of $[m]$ such that $\lambda(i)+\mu(i)+w_{i}=e\quad\text{for all }1\leq i\leq m\,$?
@@
Consider the instance of RN3DM with $W=\{1,3,4,4\}$, so $m=4$. Note that we have $\sum_{i=1}^{m}w_{i}+m(m+1)=32=m\cdot e$ for $e=8$. Let $\lambda=(4,1,3,2)$ and $\mu=(3,4,1,2)$. One can check that for every $i\in[4]$, $\lambda(i)+\mu(i)+w_{i}=e.$ For instance, $\lambda(1)+\mu(1)+w_{1}=4+3+1=8$, and $\lambda(2)+\mu(2)+w_{2}=1+4+3=8.$ The same holds for $i=3,4$. Hence $\lambda$ and $\mu$ certify that $W$ is a yes-instance of RN3DM.
Given that RN3DM is strongly NP-hard [29], we assume that the numbers in $W$ are encoded in unary. We present reductions from RN3DM to Telephone Broadcasting in cycle-star and melon graphs. For cycle-star graphs, our approach introduces an intermediate problem as part of the reduction series, whereas for melon graphs, we directly reduce RN3DM to Telephone Broadcasting.
3.1 Hardness of Telephone Broadcasting in Cycle-star Graphs
We prove that the cycle-star Telephone Broadcasting is NP-hard through a series of reductions starting from RN3DM. As a first step, we define an intermediate problem, called Even-Odd RN3DM, which is a variant of RN3DM. In this variant, instead of using two permutations $\lambda$ and $\mu$ of $[m]$ (matched with the input set $W$), we use $\alpha$ as a permutation of the even integers in $[2m]$ and $\beta$ as a permutation of the odd integers in $[2m]$ to be matched with an input set that we call $C$. Formally, Even-Odd RN3DM is defined as follows.